OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4,-2) and perpendicular to y=-x+6 A) y=-x+6 B) y=x-6 C) y=-x-2 D) y=x+2

OpenStudy (sparrow2):

if it is a perpendicular then k=1 so -2=-4+b b=2 so you have y=x+2 d

OpenStudy (anonymous):

Ok! Thanks so much!:)

OpenStudy (triciaal):